# Low-Degree Polynomials Are Good Extractors

### v1

@@ 4.2 Low-degree polynomials extract from small families
First, we consider the scenario of locally-samplable sources [DW12, Vio14, ACG+22]. A source $\mathbf{X}\sim\mathbb{F}_{2}^{n}$ is said to be $\ell$-local if $\mathbf{X}=g(\mathbf{U}_{m})$, where $g:\mathbb{F}_{2}^{m}\to\mathbb{F}_{2}^{n}$ is some function such that each output bit depends on at most $\ell$ input bits and $\mathbf{U}_{m}$ denotes the uniform distribution over $\mathbb{F}_{2}^{m}$. We can take $m\leq n\ell$ without loss of generality, in which case there are at most
$\left(\binom{n\ell}{\ell}2^{2^{\ell}}\right)^{n}\leq(en)^{\ell n}\cdot 2^{2^{\ell}\cdot n}\leq 2^{2\ell n\log n}\cdot 2^{2^{\ell}\cdot n}=2^{n(2\ell\log n+2^{\ell})}.$
$\ell$-local sources of length $n$, provided that $n\geq 4$. If $k=Cdn^{1/d}(\ell\log n+2^{\ell})^{1/d}$, we have that
$\binom{k}{\leq d}\geq\binom{k}{d}\geq(k/d)^{d}=C^{d}n(\ell\log n+2^{\ell}).$
Let $c$ be the absolute constant from Theorem 1. If we take $C$ to be large enough depending only on $c$ so that $c\cdot C^{d}n(\ell\log n+2^{\ell})>n(2\ell\log n+2^{\ell})$, we conclude that there are fewer than $2^{c\binom{k}{\leq d}}$ sources in the family. Combining a union bound over these sources with Theorem 1 immediately yields the following result.
@@
[…]
Corollary 2 both improves on and simplifies the proof of [ACG+22, Theorem 1.1], which required min-entropy $k\geq C2^{\ell}\ell^{2}d(2^{\ell}n\log n)^{1/d}$ and error $\varepsilon=2^{-\frac{ck}{d^{3}2^{\ell}\ell^{2}}}$ for some absolute constants $c,C>0$ and an intricate initial reduction to local non-oblivious bit-fixing sources. In particular, observe that when $d>\ell$, we now get that the min-entropy is $O(d(n\log n)^{1/d})$. Previously, this only happened when $d>2^{\ell}$.
[…]
@@
A random variable $\mathbf{X}\sim\mathbb{F}_{2}^{n}$ is a degree-$r$ polynomial source if there exist $\mathbb{F}_{2}$-polynomials $p_{1},\dots,p_{n}\in\mathbb{F}_{2}[x_{1},\dots,x_{m}]$ of degree at most $r$ for some positive integer $m$ (the input length) such that $\mathbf{X}=P(\mathbf{U}_{m})$, where $P=(p_{1},\dots,p_{n})$ and $\mathbf{U}_{m}$ is uniform over $\mathbb{F}_{2}^{m}$. This is an very challenging model to extract from, and several papers have attempted to do so [DGW09, BG13, GVJZ23, CGG24]. In the most recent work, Chattopadhyay, Goodman, and Gurumukhani [CGG24] established the following input reduction lemma.
@@
A random variable $\mathbf{X}\sim\mathbb{F}_{2}^{n}$ is a variety source of degree $r$ and min-entropy $k$ if for some $t$ there exist polynomials $p_{1},\dots,p_{t}\in\mathbb{F}_{2}[x_{1},\dots,x_{n}]$ of degree at most $r$ such that $\mathbf{X}$ is uniformly distributed over the variety
@@ 4.3 Low-degree polynomials are good linear seeded extractors
In this section we study the parameters of strong linear seeded extractors computable by low-degree polynomials. Such objects are crucial in the construction of several affine extractors [Li11, LZ23]. We start with the following basic definitions.
@@
We say that a string $x\in\mathbb{F}_{2}^{n}$ is $\varepsilon$-balanced if it has Hamming weight $\frac{1-\varepsilon}{2}n\leq\|x\|_{0}\leq\frac{1+\varepsilon}{2}n$. We say that a code $C\subseteq\mathbb{F}_{2}^{n}$ is $\varepsilon$-balanced [^12] Another standard name for this notion is “$\varepsilon$-biased code”. if all $x\in C$ are $\varepsilon$-balanced. We say that a code $C\subseteq\mathbb{F}_{2}^{n}$ is $\delta$-almost $\varepsilon$-balanced if $\Pr_{x\sim C}[x\text{ is not $\varepsilon$-balanced}]\leq\delta$.
@@
$\mathsf{RM}(n,d)=\{(f(x))_{x\in\mathbb{F}_{2}^{n}}\mid f\in\mathbb{F}_{2}[x_{1},\dots,x_{n}],\deg f\leq d\}.$
We will need to import the following results. The first corollary about balancedness of $\mathsf{RM}(n,d)$ follows directly from [BHL12, Lemma 1.2], which corresponds to setting $k=n$ in Theorem 1.
@@
Every linear $\varepsilon$-balanced code of length $T$ has list-size at most $2T$ at radius $\frac{1-\sqrt{\varepsilon}}{2}$. [^13] A code $C\subseteq\mathbb{F}_{q}^{n}$ has list-size at most $L$ at radius $\rho$ if |$|\{c\in C:\|c-x\|_{0}\leq\rho\}|\leq L$ for all vectors $x\in\mathbb{F}_{q}^{n}$, where $\|\cdot\|_{0}$ denotes Hamming weight.
@@
Let $G\in\mathbb{F}_{2}^{T\times n}$ be the generator matrix of a linear code with list size at most $L$ at radius $\leq\frac{1-\varepsilon}{2}$, and let $t=\log T$. Then the function $\mathsf{Ext}_{G}:\mathbb{F}_{2}^{n}\times\mathbb{F}_{2}^{t}\to\mathbb{F}_{2}$ defined as $\mathsf{Ext}_{G}(x,y):=(Gx)_{y}=\langle x,G_{y}\rangle$ is a $(k,\varepsilon)$-strong seeded extractor for min-entropy $k=\log L+\log(1/\varepsilon)+1$.
@@
The claim that $\mathsf{Ext}_{GH}$ is a strong seeded extractor follows immediately by combining the fact that the Reed-Muller code $\mathsf{RM}(t,d)$ is linear and $\delta$-almost $\varepsilon$-balanced (Corollary 5) with the fact that random subcodes of linear $\delta$-almost $\varepsilon$-balanced codes give strong seeded extractors (Theorem 4). All that remains is to show that $\mathsf{Ext}_{GH}$ will have left-degree at most $1$ and right-degree at most $d$ for any $H$. To see why, recall that, since $G\in\mathbb{F}_{2}^{T\times\binom{t}{\leq d}}$ is a generator of $\mathsf{RM}(t,d)$, there exist degree $\leq d$ polynomials $p_{1},\dots,p_{\binom{t}{\leq d}}\in\mathbb{F}_{2}[x_{1},\dots,x_{t}]$ whose truth tables make up the columns of $G$. Thus, if we let $L_{i}:\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}$ define the linear function $L_{i}(x):=\langle H_{i},x\rangle$ for each row $H_{i}$ of $H$, we can write $\mathsf{Ext}(x,y)$ as
@@ 5.1 Low-degree polynomials disperse from sumset sources
As a warmup, in this section we prove the disperser version of Theorem 2.
@@
There exist absolute constants $c,C>0$ such that for any positive integers $n$, $d\leq\frac{c\log n}{\log\log n}$, and min-entropy $k\geq Cdn^{2/d}$, a random degree $\leq d$ polynomial $f\in\mathbb{F}_{2}[x_{1},\dots,x_{n}]$ is a $k$-sumset disperser with probability at least $1-2^{-c\binom{k}{\leq d}}$.
@@
There exists an absolute constant $C>0$ such that the following holds for all positive integers $n\geq k\geq C(1+\log n)$: Let $A,B\subseteq\mathbb{F}_{2}^{n}$ be sets of size $2^{k}$. Then, there exist subsets $A^{\prime}\subseteq A$ and $B^{\prime}\subseteq B$ such that $|A^{\prime}|,|B^{\prime}|=\binom{k-C\log n}{\leq d/2}$ and $\operatorname{rank}(\operatorname{eval}_{d}(A^{\prime}+B^{\prime}))\geq\binom{k-C\log n}{\leq d}$.
Before we proceed to the (simple) proof of Lemma 14, we note some basic properties of evaluation vectors. First, it is well known that $\operatorname{rank}(\operatorname{eval}_{d}(\mathcal{B}^{n}_{d}(0)))=\binom{n}{\leq d}$, as $\mathcal{B}^{n}_{d}(0)$ is an interpolating set for degree $\leq d$ polynomials. Moreover, we have the following property.
Lemma 15.
@@
where we have invoked Lemma 15. ∎
Since it will prove useful in Section 6, we state here an alternative version of Lemma 14 which is more useful when $k$ is very small. Its proof follows in exactly the same manner as that of Lemma 14, with the exception that we invoke Lemma 4 instead of Lemma 5.
@@
The following holds for all positive integers $n\geq k>100$: Let $A,B\subseteq\mathbb{F}_{2}^{n}$ be sets of size $2^{k}$. Then, there exist subsets $A^{\prime}\subseteq A$ and $B^{\prime}\subseteq B$ such that $|A^{\prime}|,|B^{\prime}|=\binom{k/100}{\leq d/2}$ and $\operatorname{rank}(\operatorname{eval}_{d}(A^{\prime}+B^{\prime}))\geq\binom{k/100}{\leq d}$.
@@
If $d\leq\frac{c_{1}\log n}{\log\log n}$ and $k\geq c_{2}(1+\log n)$ for some absolute constants $c_{1},c_{2}>0$, we have that $\binom{t}{\leq d}\geq c_{3}\binom{k}{\leq d}$ for some absolute constant $c_{3}>0$. Plugging this into Equation 10, we conclude that the failure probability is at most
$2^{2n\binom{k}{\leq d/2}+1}\cdot 2^{-c_{3}\binom{k}{\leq d}}.$ (11)
If $k=Cdn^{2/d}$, we have
$2n\binom{k}{\leq d/2}\leq dn\left(\frac{2ek}{d}\right)^{d/2}=dn^{2}\cdot(2eC)^{d/2}.$
On the other hand, we also have that
$\binom{k}{\leq d}\geq(k/d)^{d}=C^{d}n^{2}.$
Therefore, taking $C>0$ to be sufficiently large depending only on $c_{3}$, we get that $2n\binom{k}{\leq d/2}\leq\frac{c_{3}}{2}\binom{k}{\leq d}$. Plugging this into Equation 11, we conclude that the probability that $f$ is constant on $A+B$ is at most
$2^{2n\binom{k}{\leq d/2}+1}\cdot 2^{-c_{3}\binom{k}{\leq d}}\leq 2^{-\frac{c_{3}}{2}\binom{k}{\leq d}}.$
Setting $c=c_{3}/2$ concludes the proof. ∎
@@ 5.2 Low-degree polynomials extract from sumset sources
In this section we prove the extractor part of Theorem 2. We begin with the following definition of a special type of sumset sources, for which it will be easy to show that random low-degree polynomials extract.
@@
For any $\varepsilon>0$ and positive integers $n$, $k$, and $d$ such that $n\geq k\geq 48d(n/\varepsilon^{2})^{2/d}$, the following holds with probability at least $1-2^{-\binom{k/6}{d/2}}$ for a random degree $d$ polynomial $f\in\mathbb{F}_{2}[x_{1},\dots,x_{n}]$. For any sets $\mathcal{X},\mathcal{Y}\subseteq\mathbb{F}_{2}^{n}$ of size $t\geq\binom{k/6}{d/2}$ satisfying $\operatorname{rank}(\operatorname{eval}_{d}(\mathcal{X}+\mathcal{Y}))=t^{2}$, we have that
@@
where $\mathbf{X}$ and $\mathbf{Y}$ are uniformly distributed over $\mathcal{X}$ and $\mathcal{Y}$, respectively.
@@
Fix any sets $\mathcal{X},\mathcal{Y}\subseteq\mathbb{F}_{2}^{n}$ of size $t\geq\binom{k/6}{d/2}$ satisfying $\operatorname{rank}(\operatorname{eval}_{d}(\mathcal{X}+\mathcal{Y}))=t^{2}$. Then, it follows that the $t^{2}$ random variables $(f(x+y))_{(x,y)\in\mathcal{X}\times\mathcal{Y}}$ are independent and uniformly distributed over $\mathbb{F}_{2}$. Since
$\operatorname{bias}_{\mathbf{X}+\mathbf{Y}}(f)=\mathbb{E}_{(x,y)\sim\mathcal{X}\times\mathcal{Y}}[(-1)^{f(x+y)}]=\frac{1}{t^{2}}\sum_{(x,y)\in\mathcal{X}\times\mathcal{Y}}(-1)^{f(x+y)}$
and $\mathbb{E}_{f}[(-1)^{f(x+y)}]=0$, a direct application of the Chernoff bound (Lemma 1) ensures that
$\Pr_{f}[|\operatorname{bias}_{\mathbf{X}+\mathbf{Y}}(f)|>\varepsilon]\leq 2\cdot 2^{-\varepsilon^{2}t^{2}/4}.$
By taking a union bound over all the at most $\binom{2^{n}}{t}^{2}\leq 2^{2nt}$ choices of $\mathcal{X}$ and $\mathcal{Y}$, we conclude that the probability that $|\operatorname{bias}_{\mathbf{X}+\mathbf{Y}}(f)|>\varepsilon$ for some sets $\mathcal{X}$ and $\mathcal{Y}$ is at most
$2^{2nt}\cdot 2\cdot 2^{-2\varepsilon^{2}t^{2}}=2^{1+2t(n-\varepsilon^{2}t/8)}.$ (12)
Recall that $k\geq 48d(n/\varepsilon^{2})^{2/d}$. The quantity in Equation 12 will be at most $2^{1-2t}\leq 2^{1-2\binom{k/6}{d/2}}\leq 2^{-\binom{k/6}{d/2}}$ provided that $t\geq\frac{8(n+1)}{\varepsilon^{2}}$. This holds since
$t\geq\binom{k/6}{d/2}\geq\left(\frac{k}{3d}\right)^{d/2}\geq 16n/\varepsilon^{2}\geq 8(n+1)/\varepsilon^{2}.\qed$
@@
There exists a constant $c>0$ such that the following holds for all positive integers $n\geq k$ and $d\leq k/3$. Let $\mathbf{W}=\mathbf{X}+\mathbf{Y}$ be an $(n,k)$-sumset source. Then, $\mathbf{W}$ is $2^{-ck}$-close to a convex combination of flat sumset sources $\mathbf{W}^{\prime}=\mathbf{X}^{\prime}+\mathbf{Y}^{\prime}$ with full evald-rank and such that $|\mathsf{supp}(\mathbf{X}^{\prime})|,|\mathsf{supp}(\mathbf{Y}^{\prime})|\geq\binom{k/6}{d/2}$.
@@
$\mathcal{B}_{0}=\{u\in\mathbb{F}_{2}^{m}:w(u)=d/2,\mathsf{supp}(u)\subseteq\{1,\dots,k/3\}\},$
$\mathcal{B}_{1}=\{v\in\mathbb{F}_{2}^{m}:w(v)=d/2,\mathsf{supp}(v)\subseteq\{2k/3+1,\dots,k\}\}.$

### v2

@@ 4.2 Low-degree polynomials extract from small families
First, we consider the scenario of locally-samplable sources [DW12, Vio14, ACG+22]. A source $\mathbf{X}\sim\mathbb{F}_{2}^{n}$ is said to be $r$-local if $\mathbf{X}=g(\mathbf{U}_{m})$, where $g:\mathbb{F}_{2}^{m}\to\mathbb{F}_{2}^{n}$ is some function such that each output bit depends on at most $r$ input bits and $\mathbf{U}_{m}$ denotes the uniform distribution over $\mathbb{F}_{2}^{m}$. We can take $m\leq nr$ without loss of generality, in which case there are at most
$\left(\binom{nr}{r}2^{2^{r}}\right)^{n}\leq(en)^{rn}\cdot 2^{2^{r}\cdot n}\leq 2^{2rn\log n}\cdot 2^{2^{r}\cdot n}=2^{n(2r\log n+2^{r})}$
$r$-local sources of length $n$, provided that $n\geq 4$. If $k=Cdn^{1/d}(r\log n+2^{r})^{1/d}$, we have that
$\binom{k}{\leq d}\geq\binom{k}{d}\geq(k/d)^{d}=C^{d}n(r\log n+2^{r}).$
Let $c$ be the absolute constant from Theorem 1. If we take $C$ to be large enough depending only on $c$ so that $c\cdot C^{d}n(r\log n+2^{r})>n(2r\log n+2^{r})$, we conclude that there are fewer than $2^{c\binom{k}{\leq d}}$ sources in the family. Combining a union bound over these sources with Theorem 1 immediately yields the following result.
@@
For every $\delta\in(0,1)$ there exist constants $c,C>0$ such that the following holds for all $n$: For $k\geq Cdn^{1/d}(r\log n+2^{r})^{1/d}$ and $d\leq(1-\delta)k$, a random degree $d$ polynomial is an $(\varepsilon=2^{-ck/d})$-extractor for the family of length-$n$ $r$-local sources of min-entropy $k$ with probability at least $1-2^{-c\binom{k}{\leq d}}$ over the choice of $f$.
Corollary 2 both improves on and simplifies the proof of [ACG+22, Theorem 1.1], which required min-entropy $k\geq C2^{r}r^{2}d(2^{r}n\log n)^{1/d}$ and error $\varepsilon=2^{-\frac{ck}{d^{3}2^{r}r^{2}}}$ for some absolute constants $c,C>0$ and an intricate initial reduction to local non-oblivious bit-fixing sources. In particular, observe that when $d>r$, we now get that the min-entropy is $O(d(n\log n)^{1/d})$. Previously, this only happened when $d>2^{r}$.
It is also instructive to compare our improved result with the lower bound from [ACG+22, Theorem 1.2], which states that no degree-$d$ polynomial extracts from length-$n$ $r$-local sources of min-entropy $k=cd(rn\log n)^{1/d}$ for some absolute constant $c>0$. For example, when the locality satisfies $r<\log\log n$, Corollary 2 is optimal up to the constant factor.
@@
A random variable $\mathbf{X}\sim\mathbb{F}_{2}^{n}$ is a degree-$r$ polynomial source if there exist $\mathbb{F}_{2}$-polynomials $p_{1},\dots,p_{n}:\mathbb{F}_{2}^{m}\to\mathbb{F}_{2}$ of degree at most $r$ for some positive integer $m$ (the input length) such that $\mathbf{X}=P(\mathbf{U}_{m})$, where $P=(p_{1},\dots,p_{n})$ and $\mathbf{U}_{m}$ is uniform over $\mathbb{F}_{2}^{m}$. This is a very challenging model to extract from, and several papers have attempted to do so [DGW09, BG13, GVJZ23, CGG24]. In the most recent work, Chattopadhyay, Goodman, and Gurumukhani [CGG24] established the following input reduction lemma.
@@
A random variable $\mathbf{X}\sim\mathbb{F}_{2}^{n}$ is a variety source of degree $r$ and min-entropy $k$ if for some $t$ there exist polynomials $p_{1},\dots,p_{t}:\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}$ of degree at most $r$ such that $\mathbf{X}$ is uniformly distributed over the variety
@@ 4.3 Low-degree polynomials are good linear seeded extractors
In this section we study the parameters of strong linear seeded extractors computable by low-degree polynomials. Such objects are crucial in the construction of several affine extractors [Li11, LZ24]. We start with the following basic definitions.
@@
We say that a string $x\in\mathbb{F}_{2}^{n}$ is $\varepsilon$-balanced if it has Hamming weight $\frac{1-\varepsilon}{2}n\leq\mathsf{wt}(x)\leq\frac{1+\varepsilon}{2}n$. We say that a code $C\subseteq\mathbb{F}_{2}^{n}$ is $\varepsilon$-balanced [^14] Another standard name for this notion is “$\varepsilon$-biased code.” if all nonzero $x\in C$ are $\varepsilon$-balanced. We say that a code $C\subseteq\mathbb{F}_{2}^{n}$ is $\delta$-almost $\varepsilon$-balanced if $\Pr_{x\sim C}[x\text{ is not $\varepsilon$-balanced}]\leq\delta$.
@@
$\mathsf{RM}(n,d)=\{(f(x))_{x\in\mathbb{F}_{2}^{n}}\mid f:\mathbb{F}_{2}^{n}\to\mathbb{F}_{2},\deg f\leq d\}.$
We will need to import the following results. The first corollary about balancedness of $\mathsf{RM}(n,d)$ follows directly from [BHL12, Lemma 1.2] (Theorem ), which corresponds to setting $k=n$ in Theorem 1.
@@
Every linear $\varepsilon$-balanced code of length $T$ has list-size at most $2T$ at radius $\frac{1-\sqrt{\varepsilon}}{2}$. [^15] A code $C\subseteq\mathbb{F}_{q}^{n}$ has list-size at most $L$ at radius $\rho$ if |$|\{c\in C:\mathsf{wt}(c-x)\leq\rho n\}|\leq L$ for all vectors $x\in\mathbb{F}_{q}^{n}$.
@@
Let $G\in\mathbb{F}_{2}^{T\times n}$ be the generator matrix of a linear code with list size $\leq L$ at radius $\leq\frac{1-\varepsilon}{2}$, and let $t=\log T$. Then the function $\mathsf{Ext}_{G}:\mathbb{F}_{2}^{n}\times\mathbb{F}_{2}^{t}\to\mathbb{F}_{2}$ defined as $\mathsf{Ext}_{G}(x,y):=(Gx)_{y}=\langle x,G_{y}\rangle$ is a $(k,\varepsilon)$-strong seeded extractor for min-entropy $k=\log L+\log(1/\varepsilon)+1$.
@@
The claim that $\mathsf{Ext}_{GH}$ is a strong seeded extractor follows immediately by combining the fact that the Reed-Muller code $\mathsf{RM}(t,d)$ is linear and $\delta$-almost $\varepsilon$-balanced (Corollary 5) with the fact that random subcodes of linear $\delta$-almost $\varepsilon$-balanced codes give strong seeded extractors (Theorem 3). All that remains is to show that $\mathsf{Ext}_{GH}$ will have left-degree at most $1$ and right-degree at most $d$ for any $H$. To see why, recall that, since $G\in\mathbb{F}_{2}^{T\times\binom{t}{\leq d}}$ is a generator of $\mathsf{RM}(t,d)$, there exist degree $\leq d$ polynomials $p_{1},\dots,p_{\binom{t}{\leq d}}:\mathbb{F}_{2}^{t}\to\mathbb{F}_{2}$ whose truth tables make up the columns of $G$. Thus, if we let $L_{i}:\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}$ define the linear function $L_{i}(x):=\langle H_{i},x\rangle$ for each row $H_{i}$ of $H$, we can write $\mathsf{Ext}(x,y)$ as
@@ 5.1 Low-degree polynomials disperse from sumset sources
As a warmup, we prove a disperser version of Theorem 2 (with a slightly better bound on the probability). For simplicity, we focus here on the case of even degree $d$.
@@
There exist constants $C,c>0$ such that for any $n\geq k\geq d\in\mathbb{N}$ (with $d$ even) satisfying $d\leq\frac{c\log n}{\log\log n}$ and $k\geq Cdn^{2/d}$, a random degree $d$ polynomial $f:\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}$ is a disperser for $(n,k)$-sumset sources, with probability at least $1-2^{-c\binom{k}{\leq d}}$.
@@
There exists a constant $C>0$ such that for all $n\geq k\geq d\in\mathbb{N}$ (with $d$ even) satisfying $k\geq C(1+\log n)$, the following holds. Let $A,B\subseteq\mathbb{F}_{2}^{n}$ be sets of size $2^{k}$. Then, there exist subsets $A^{\prime}\subseteq A$ and $B^{\prime}\subseteq B$ such that $|A^{\prime}|,|B^{\prime}|=\binom{k-C\log n}{\leq d/2}$ and $\operatorname{rank}(\operatorname{eval}_{d}(A^{\prime}+B^{\prime}))\geq\binom{k-C\log n}{\leq d}$.
Before we proceed to the (simple) proof of Lemma 12, we note some basic properties of evaluation vectors. First, it is well known that $\operatorname{rank}(\operatorname{eval}_{d}(\mathcal{B}^{n}_{d}(0)))=\binom{n}{\leq d}$, as $\mathcal{B}^{n}_{d}(0)$ is an interpolating set for degree $\leq d$ polynomials (see, e.g., [O’D14, Proposition 6.21]). Moreover, we have the following property.
Claim 2.
@@
By combining the above property with the leftover hash lemma, we can now prove Lemma 12.
@@
where we have invoked Claim 2. ∎
Next, since it will be useful in Section 6, we state here an alternative version of Lemma 12, which is more helpful when $k$ is very small. Its proof follows in exactly the same manner as that of Lemma 12, with the exception that we invoke Lemma 4 instead of Lemma 5.
@@
For all $n\geq k\geq d\in\mathbb{N}$ (with $d$ even) satisfying $k>100$, the following holds. Let $A,B\subseteq\mathbb{F}_{2}^{n}$ be sets of size $2^{k}$. Then, there exist subsets $A^{\prime}\subseteq A$ and $B^{\prime}\subseteq B$ such that $|A^{\prime}|,|B^{\prime}|=\binom{k/100}{\leq d/2}$ and $\operatorname{rank}(\operatorname{eval}_{d}(A^{\prime}+B^{\prime}))\geq\binom{k/100}{\leq d}$.
@@
We will now invoke 1 with $\beta=\delta=1/2$, $k$ in place of $n$, and $t=k-C_{0}(1+\log n)$. Let $\gamma>0$ be the absolute constant correspoding to $\beta=\delta=1/2$ in 1. By hypothesis, we have $k\geq Cdn^{2/d}$ and $d\leq\frac{c\log n}{\log\log n}<k/2=\delta k$ for $C>0$ an appropriately large constant and $c>0$ an appropriately small constant. In particular, this means that we can enforce $10<c<$ 1 and choose $C>0$ large enough depending on $\gamma$ and $C_{0}$ so that $k\geq Cdn^{\frac{2\log\log n}{\log n}}\geq Cd\log n$ satisfies $t=k-C_{0}(1+\log n)\geq d$ and $t\geq(1-\gamma/d)k$. By 1, we then get that $\binom{t}{\leq d}\geq\frac{1}{2}\binom{k}{\leq d}$. Plugging this into Equation 10, we conclude that the failure probability is at most
$2^{2n\binom{k}{\leq d/2}+1}\cdot 2^{-\frac{1}{2}\binom{k}{\leq d}}.$ (11)
We also have
$2n\binom{k}{\leq d/2}\leq 2dn\left(\frac{2ek}{d}\right)^{d/2}\leq\frac{1}{4}(k/d)^{d}\leq\frac{1}{4}\binom{k}{\leq d},$
where the second inequality uses that $k\geq Cdn^{2/d}$ for a sufficiently large constant $C>0$. Plugging this into Equation 11, we conclude that the probability that $f$ is constant on $A+B$ is at most
$2^{2n\binom{k}{\leq d/2}+1}\cdot 2^{-\frac{1}{2}\binom{k}{\leq d}}\leq 2^{-\frac{1}{4}\binom{k}{\leq d}+1}.\qed$
@@ 5.2 Low-degree polynomials extract from sumset sources
In this section, we prove Theorem 2 in full generality. We begin with the following definition of a special type of sumset sources, for which it will be easy to show that random low-degree polynomials extract.
@@
For any $t,n\in\mathbb{N}$ and $\varepsilon>0$ such that $t\geq 64n/\varepsilon^{2}$, the following holds with probability at least $1-2\cdot 2^{-\varepsilon^{2}t^{2}/32}$ over the selection of a random degree $d$ polynomial $f:\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}$. For any sets $X,Y\subseteq\mathbb{F}_{2}^{n}$ of size $t$ satisfying $\operatorname{rank}(\operatorname{eval}_{d}(X+Y))=t^{2}$, we have that
@@
where $\mathbf{X}$ and $\mathbf{Y}$ are uniformly distributed over $X$ and $Y$, respectively.
@@
Fix any sets $X,Y\subseteq\mathbb{F}_{2}^{n}$ of size $t$ satisfying $\operatorname{rank}(\operatorname{eval}_{d}(X+Y))=t^{2}$. By this rank condition, we know that the $t^{2}$ random variables $\{f(x+y)\}_{(x,y)\in X\times Y}$ are independent and uniform over $\mathbb{F}_{2}$. Thus, if we define for every $(x,y)\in X\times Y$ the indicator random variable $\mathbf{Z}_{x,y}:=1[f(x+y)=0]$, then
$\operatorname{bias}_{\mathbf{X}+\mathbf{Y}}(f)=2\left(\frac{1}{t^{2}}\sum_{x,y}\mathbf{Z}_{x,y}-\frac{1}{2}\right)=2\left(\frac{1}{t^{2}}\sum_{x,y}\mathbf{Z}_{x,y}-\mathbb{E}\left[\frac{1}{t^{2}}\sum_{x,y}\mathbf{Z}_{x,y}\right]\right),$
and a direct application of the Chernoff bound (Lemma 1) yields
$\Pr_{f}\left[|\operatorname{bias}_{\mathbf{X}+\mathbf{Y}}|>\varepsilon\right]\leq 2e^{-\varepsilon^{2}t^{2}/16}.$
Finally, by taking a union bound over all the at most $\binom{2^{n}}{t}^{2}\leq 2^{2nt}$ choices of $X$ and $Y$, we conclude that the probability that $|\operatorname{bias}_{\mathbf{X}+\mathbf{Y}}(f)|>\varepsilon$ for some sets $X,Y$ is at most
$2^{2nt}\cdot 2e^{-\varepsilon^{2}t^{2}/16},$
which is at most $2\cdot 2^{-\varepsilon^{2}t^{2}/32}$ when $t\geq 64n/\varepsilon^{2}$, as desired. ∎
@@
There exists a constant $c>0$ such that for all $n\geq k\geq d\in\mathbb{N}$, the following holds. Let $\mathbf{W}=\mathbf{X}+\mathbf{Y}$ be an $(n,k)$-sumset source. Then $\mathbf{W}$ is $2^{-ck}$-close to a convex combination of flat sumset sources $\mathbf{W}^{\star}=\mathbf{X}^{\star}+\mathbf{Y}^{\star}$ with full evald-rank and such that $|\mathsf{supp}(\mathbf{X}^{\star})|=|\mathsf{supp}(\mathbf{Y}^{\star})|=\binom{k/6}{\lfloor d/2\rfloor}$.
@@
$\mathcal{B}_{0}=\{u\in\mathbb{F}_{2}^{m}:\mathsf{wt}(u)=\lfloor d/2\rfloor,\mathsf{supp}(u)\subseteq\{1,\dots,m/3\}\},$
$\mathcal{B}_{1}=\{v\in\mathbb{F}_{2}^{m}:\mathsf{wt}(v)=\lfloor d/2\rfloor,\mathsf{supp}(v)\subseteq\{2m/3+1,\dots,m\}\}.$